What is the probability that the sum of the dice is odd and both dice show the number 3?
The probability is 0. If both dice show the number 3 then the sum is 6 which is not odd.
How many ways can 9 books be arranged on a shelf if there is only room for 5 of the books?
9 x 8 x 7 x 6 x 5
It is 84.3%
You randomly select one card from a 52-card deck. Find the probability of selecting the king of diamonds or the jack of
The total probability of something happening plus the probability of that same thing not happening is 1, or 100 %
→ probability of not happening = 1 - 0.25 = 0.75 or 100 % - 25 % = 75 %
What is the probability of choosing a prime number?
Although there are infinitely many primes, they become rarer and rarer so that as the number of numbers increases, the probability that picking one of them at random is a prime number tends to zero*.
In the first 10 numbers there are 4 primes, so the probability of picking one is 4/10 = 2/5 = 0.4
In the first 100 numbers there are 26 primes, so the probability of picking one is 25/100 = 1/4 = 0.25
In the first 1,000 numbers there are 169 primes, so the probability of picking one is 168/1000 = 0.168
In the first 10,000 numbers there are 1,229 primes, so the probability of picking one is 0.1229
In the first 100,000 numbers there are 9592 primes, so the probability of picking one is 0.09592
In the first 1,000,000 numbers there are 78,498 primes, so the probability of picking one is 0.078498
In the first 10,000,000 numbers there are 664,579 primes, so the probability of picking one is 0.0664579
* Given any small value ε less than 1 and greater than 0, it is possible to find a number n such that the probability of picking a prime at random from the numbers 1-n is less than the given small value ε.
There are no graphs to choose from!
What is the probability that both cards are spades?
The answer will depend on:
It also depends on how many times the experiment - of drawing two cards - is repeated. If repeated a sufficient number of times the probability will be so close to 1 as to make no difference from a certainty.
It is 0.25
When h has the value 5 what is 6h 9?
Unfortunately, limitations of the browser used by Answers.com means that we cannot see most symbols. It is therefore impossible to give a proper answer to your question. Please resubmit your question spelling out the symbols as "plus", "minus", "times", "divided by", "equals", "squared", "cubed" etc. There is no symbol between 6h and 9.
When you have 2 six sided dice what is the probability that the largest number rolled is a 3?
It is 1/6.
There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.
If you believe that the children's genders are not independent then you would need to get empirical evidence from all families with four or more children in which the first three children were girls. If there are g families in which the fourth is a girl and b where the fourth is a boy then the required probability is b/(g+b).
However, if you assume that the children's genders are independent events then, given that the probability of a boy is approx 0.52, the probability of the fourth child is a boy is 0.52
"Playing cards" are chosen at random.
How many solutions are there to the equation , where , i = 1, 2, 3, 4, 5, is a nonnegative integer such that
A die has six sides, so the probability of rolling an even number is 1 in 2, or 50-50. A coin has two sides, so the probability of flipping the coin and getting heads is 1 in 2, or 50-50.
The probability that both will happen together is the one in two OF one in two, or one in FOUR chance that both will happen.
So, the probability is 25%.
If this is the only information that you have then you must use the Poisson distribution.